OpenStudy (anonymous):

how do you find the x-intercept of y = 4x - 64 *ill take an explanation before an answer, thanks

OpenStudy (pfenn1):

An x-intercept is a point on the graph where y is zero. so plug y into the equation and solve for x.

OpenStudy (pfenn1):

Sorry. An x-intercept is a point on the graph where y is zero. So plug 0 for y into the equation and solve for x.

OpenStudy (anonymous):

when i plugged in 0 for y i got 16 which wasnt my answer choice

OpenStudy (pfenn1):

Weird. That is what I got also. Please double check that you copied the equation correctly and that you are looking of the x-intercept and not the y-intercept.

OpenStudy (pfenn1):

The equation is \[y=4^x-64\]

OpenStudy (anonymous):

oh i see, the x is raised

OpenStudy (anonymous):

do you know this kind of problem?

OpenStudy (pfenn1):

I think you need to use log. \[0=4^x-64\rightarrow4^x = 64\]

OpenStudy (anonymous):

so that'l be A. (3,0)

OpenStudy (pfenn1):

I don't remember how to solve this but I did notice that 4^2=16 and 4^3=64 so that x=3.

OpenStudy (pfenn1):

Yes, you got it.

OpenStudy (anonymous):

Thanks for your help

OpenStudy (pfenn1):

You're welcome.